A proton, traveling with a velocity of 6.8 × 106 m/s due east, experiences a maximum magnetic force of 8.0 × 10-14 N. The direction of the force is straight down, toward the surface of the earth. What is the magnitude and direction of the magnetic field B, assumed perpendicular to the motion?


1
Expert's answer
2021-03-17T16:56:00-0400

We can find the magnitude of the magnetic field as follows:


"F_B=qvBsin\theta,"

"B=\dfrac{F_B}{qvsin\theta}=\dfrac{8.0\cdot10^{-14}\ N}{1.6\cdot10^{-19}\ C\cdot6.8\cdot10^6\ \dfrac{m}{s}\cdot sin90^{\circ}}=0.073\ T."

According to the right-hand rule, the magnetic field directed upward.